Calculate the de Broglie wavelength associated with a ball of mass 0.50 kg moving
with a velocity of 20 m s−1
.
1
Expert's answer
2020-04-06T11:46:06-0400

De Broglie wavelength l = h/(m*v), where h is Plank’s constant 6.626*10^(-34) kg*m^2/s. For the given ball the de Broglie wavelength is:

l = 6.626*10^(-34)/(0.5*20) = 6.626*10^(-35) m